# TabMGP: Martingale Posterior with TabPFN

### v1

@@ Appendix H Coverage rate of the marginal credible interval
Table 7: Coverage and Winkler score (median over 100 repetitions, shown in parantheses) of the marginal credible intervals for logistic regression with synthetic data. The method with the lowest Winkler score is shown in bold. The target coverage is 0.95.
$z_{1:n}$ Feature Name TabMGP $B$ Copula Bayes
concrete intercept 0.91 (0.29) 0.92 (0.26) 1.00 (0.55) 1.00 (0.40)
cement 0.90 (0.33) 0.91 (0.30) 0.98 (0.50) 1.00 (0.48)
fly_ash 0.96 (0.34) 0.92 (0.33) 0.99 (0.63) 0.99 (0.50)
superplasticizer 0.91 (0.37) 0.85 (0.36) 0.99 (0.49) 0.99 (0.51)
coarse_aggregate 0.91 (0.30) 0.91 (0.28) 0.99 (0.47) 0.99 (0.43)
fine_aggregate 0.95 (0.29) 0.95 (0.27) 1.00 (0.45) 1.00 (0.44)
age 0.83 (0.36) 0.83 (0.36) 0.83 (0.34) 0.89 (0.41)
quake intercept 0.96 (0.48) 0.95 (0.51) 1.00 (1.02) 1.00 (0.56)
col_1 1.00 (0.23) 0.88 (0.48) 0.99 (0.50) 1.00 (0.63)
col_2 0.93 (0.25) 0.94 (0.52) 1.00 (0.80) 0.97 (0.57)
col_3 1.00 (0.24) 0.95 (0.50) 1.00 (0.81) 0.97 (0.58)
airfoil intercept 0.93 (0.42) 0.93 (0.37) 0.99 (0.63) 0.99 (0.57)
frequency 0.90 (0.50) 0.87 (0.46) 0.84 (0.51) 0.99 (0.63)
chord_length 0.96 (0.45) 0.91 (0.39) 0.98 (0.55) 1.00 (0.59)
free_stream_velocity 0.94 (0.40) 0.93 (0.36) 0.99 (0.55) 0.99 (0.57)
displacement_thickness 0.93 (0.46) 0.87 (0.41) 0.96 (0.59) 1.00 (0.60)
energy intercept 0.92 (0.29) 0.88 (0.15) 1.00 (0.45) 1.00 (0.58)
wall_area 1.00 (0.25) 0.96 (0.12) 1.00 (0.35) 1.00 (0.62)
overall_height 0.97 (0.26) 0.89 (0.15) 1.00 (0.49) 1.00 (0.59)
orientation 0.96 (0.25) 0.88 (0.16) 1.00 (0.34) 1.00 (0.58)
glazing_area 0.93 (0.26) 0.86 (0.16) 1.00 (0.34) 1.00 (0.60)
glazing_area_distribution 0.96 (0.25) 0.89 (0.16) 0.99 (0.33) 1.00 (0.59)
fish intercept 0.93 (0.45) 0.83 (0.37) 0.99 (0.54) 1.00 (0.58)
CIC0 0.89 (0.49) 0.86 (0.46) 0.86 (0.52) 0.99 (0.63)
SM1_Dz 0.89 (0.50) 0.83 (0.42) 0.92 (0.50) 0.98 (0.64)
GATS1i 0.92 (0.42) 0.92 (0.35) 0.96 (0.47) 1.00 (0.58)
NdsCH 0.95 (0.46) 0.81 (0.42) 0.95 (0.46) 1.00 (0.64)
NdssC 0.94 (0.49) 0.89 (0.47) 0.94 (0.49) 0.99 (0.68)
kin8nm intercept 0.94 (0.48) 0.91 (0.39) 0.90 (0.36) 1.00 (0.60)
theta1 0.96 (0.41) 0.85 (0.40) 0.86 (0.37) 0.99 (0.63)
theta2 0.98 (0.40) 0.86 (0.40) 0.88 (0.36) 1.00 (0.61)
theta3 0.96 (0.52) 0.90 (0.37) 0.78 (0.36) 0.99 (0.61)
theta4 0.96 (0.43) 0.90 (0.39) 0.90 (0.35) 1.00 (0.60)
theta5 0.82 (0.49) 0.87 (0.41) 0.86 (0.35) 0.98 (0.61)
theta6 0.92 (0.43) 0.79 (0.42) 0.78 (0.36) 0.95 (0.61)
theta7 0.94 (0.42) 0.84 (0.40) 0.84 (0.36) 0.98 (0.61)
theta8 0.96 (0.41) 0.84 (0.40) 0.87 (0.36) 0.97 (0.61)
@@
Table 8: Coverage and Winkler score (median over 100 repetitions, shown in parantheses) of the marginal credible intervals for linear regression with real data. The method with the lowest Winkler score is shown in bold. The target coverage is 0.95.
$z_{1:n}$ Feature Name TabMGP $B$ Copula Bayes
auction intercept 0.93 (0.75) 0.89 (0.55) 1.00 (1.70) 1.00 (1.24)
property.product_2 0.91 (1.41) 0.87 (1.17) 0.84 (1.71) 0.99 (1.81)
property.product_3 0.93 (1.07) 0.89 (0.88) 0.80 (1.73) 1.00 (2.08)
property.product_4 0.96 (1.06) 0.89 (0.87) 0.88 (1.51) 1.00 (2.02)
property.product_5 0.94 (1.16) 0.89 (0.91) 0.97 (1.73) 1.00 (2.32)
property.product_6 0.94 (1.00) 0.93 (0.79) 0.82 (1.61) 1.00 (1.91)
process.b1.capacity 0.98 (0.50) 0.94 (0.42) 1.00 (1.04) 1.00 (0.72)
process.b2.capacity 0.90 (0.40) 0.80 (0.35) 1.00 (0.84) 0.99 (0.64)
process.b3.capacity 0.92 (0.35) 0.94 (0.27) 0.98 (0.67) 1.00 (0.69)
process.b4.capacity 0.93 (0.40) 0.86 (0.33) 1.00 (1.20) 1.00 (0.64)
property.price 0.91 (0.46) 0.81 (0.41) 1.00 (1.03) 1.00 (0.73)
grid intercept 0.92 (0.41) 0.91 (0.31) 0.86 (0.34) 1.00 (0.63)
tau1 0.91 (0.40) 0.81 (0.31) 0.91 (0.32) 1.00 (0.63)
tau2 0.94 (0.40) 0.89 (0.31) 0.89 (0.33) 1.00 (0.62)
tau3 0.91 (0.39) 0.86 (0.31) 0.90 (0.34) 1.00 (0.62)
tau4 0.94 (0.40) 0.89 (0.31) 0.91 (0.33) 0.99 (0.63)
p2 0.97 (0.37) 0.90 (0.30) 0.96 (0.35) 1.00 (0.63)
p3 0.99 (0.36) 0.91 (0.30) 0.97 (0.34) 1.00 (0.63)
p4 0.99 (0.38) 0.87 (0.30) 0.97 (0.34) 1.00 (0.62)
g1 0.92 (0.40) 0.82 (0.32) 0.86 (0.34) 0.99 (0.63)
g2 0.98 (0.40) 0.88 (0.31) 0.94 (0.34) 1.00 (0.63)
g3 0.96 (0.40) 0.86 (0.31) 0.89 (0.34) 1.00 (0.64)
g4 0.98 (0.40) 0.92 (0.31) 0.93 (0.35) 1.00 (0.63)
abalone intercept 0.94 (1.24) 0.76 (1.23) 0.99 (1.79) 0.98 (1.76)
Sex_1 0.92 (1.83) 0.78 (1.78) 0.95 (1.88) 0.98 (2.72)
Sex_2 0.99 (1.48) 0.81 (1.70) 0.95 (1.58) 0.99 (2.31)
Height 0.99 (1.54) 0.65 (1.76) 0.96 (1.37) 0.94 (2.20)
Shucked_weight 0.96 (1.42) 0.68 (1.61) 0.92 (1.54) 0.95 (2.01)
@@
Table 9: Coverage and Winkler score (median over 100 repetitions, shown in parantheses) of the marginal credible intervals for linear regression with real data. The method with the lowest Winkler score is shown in bold. The target coverage is 0.95.
$z_{1:n}$ Feature Name TabMGP $B$ Copula Bayes
rice intercept_Osmancik 0.89 (2.72) 0.30 (14.3) 0.97 (6.22) 0.59 (3.48)
Minor_Axis_Length_Osmancik 0.93 (3.91) 0.49 (7.81) 0.97 (11.2) 0.74 (5.20)
Eccentricity_Osmancik 0.91 (5.36) 0.37 (20.1) 0.98 (14.5) 0.73 (7.45)
Extent_Osmancik 0.99 (1.21) 0.93 (2.18) 1.00 (4.14) 0.98 (2.12)
sepsis intercept_1 1.00 (1.77) 0.98 (2.91) 1.00 (6.42) 0.97 (3.24)
sex_0male_1female_1_1 1.00 (1.73) 0.99 (3.59) 1.00 (7.84) 1.00 (4.00)
age_years_1 0.85 (2.00) 0.96 (2.69) 1.00 (4.88) 0.98 (3.24)
episode_number_1 1.00 (1.19) 0.93 (2.57) 1.00 (3.28) 1.00 (3.53)
banknote intercept_2 1.00 (1.33) 1.00 (1.44) 1.00 (1.45) 1.00 (1.52)
V1_2 0.96 (2.24) 0.96 (2.63) 1.00 (2.85) 0.99 (3.06)
V2_2 0.97 (2.26) 0.98 (2.50) 0.97 (2.46) 0.99 (2.57)
V4_2 0.98 (2.01) 0.98 (2.26) 0.99 (2.31) 0.98 (2.05)
mozilla intercept_1 0.94 (1.44) 0.77 (1.77) 1.00 (2.40) 0.87 (1.58)
id_1 0.99 (0.77) 1.00 (0.91) 1.00 (1.83) 1.00 (0.96)
end_1 0.98 (1.00) 0.97 (1.11) 1.00 (2.08) 0.99 (1.08)
event_1 1.00 (0.78) 1.00 (0.92) 1.00 (2.05) 1.00 (1.00)
size_1 0.93 (4.13) 0.74 (5.06) 0.99 (4.43) 0.73 (4.33)
skin intercept_2 1.00 (1.70) 1.00 (1.85) 1.00 (2.11) 1.00 (2.89)
V1_2 1.00 (3.00) 1.00 (3.29) 1.00 (3.23) 1.00 (2.49)
V3_2 0.99 (2.91) 0.99 (3.18) 1.00 (3.37) 1.00 (2.85)
blood intercept_2 1.00 (1.48) 0.93 (2.15) 1.00 (3.68) 0.95 (2.03)
V1_2 0.82 (2.07) 0.91 (2.66) 0.98 (4.18) 0.95 (2.55)
V2_2 0.91 (2.50) 0.74 (3.73) 0.99 (4.99) 0.81 (3.38)
V4_2 0.81 (2.10) 0.79 (2.89) 1.00 (4.01) 0.87 (2.95)
phoneme intercept_2 0.97 (1.73) 0.83 (1.98) 1.00 (3.37) 0.91 (2.35)
V1_2 0.96 (1.93) 0.85 (2.15) 1.00 (3.58) 0.94 (2.97)
V2_2 0.95 (1.85) 0.87 (2.33) 0.98 (3.44) 0.93 (2.35)
V3_2 0.90 (1.93) 0.80 (2.47) 0.96 (3.70) 0.83 (2.30)
V4_2 0.91 (1.71) 0.80 (1.97) 0.95 (3.70) 0.80 (1.97)
V5_2 0.89 (1.45) 0.83 (1.95) 0.94 (3.32) 0.86 (1.76)
telescope intercept_h 0.97 (1.93) 0.88 (2.73) 1.00 (7.44) 0.94 (2.63)
fLength_h 0.87 (5.08) 0.69 (7.01) 0.98 (13.4) 0.67 (6.61)
fWidth_h 0.95 (4.47) 0.70 (6.45) 0.98 (11.9) 0.80 (6.27)
fConc1_h 0.95 (2.64) 0.76 (3.63) 0.99 (9.79) 0.81 (3.27)
fAsym_h 0.96 (2.58) 0.78 (3.42) 1.00 (7.77) 0.88 (3.67)
fM3Long_h 0.95 (2.52) 0.87 (3.81) 0.99 (7.24) 0.87 (3.71)
fM3Trans_h 0.97 (2.21) 0.88 (3.12) 1.00 (6.63) 0.90 (3.21)
fAlpha_h 0.87 (2.56) 0.79 (3.04) 0.93 (8.26) 0.75 (2.61)
fDist_h 0.93 (2.12) 0.81 (2.96) 0.98 (6.56) 0.83 (2.67)
@@
Table 10: Coverage and Winkler score (median over 100 repetitions, shown in parantheses) of the marginal credible intervals for logistic regression with real data. The method with the lowest Winkler score is shown in bold. The target coverage is 0.95.
$z_{1:n}$ Feature Name TabMGP $B$ Bayes
yeast intercept_ME2 0.89 (6.53) 0.39 (53.6) 0.41 (42.0)
intercept_ME3 0.95 (3.53) 0.54 (6.47) 0.66 (5.32)
intercept_MIT 0.97 (1.08) 0.87 (1.33) 0.92 (1.47)
intercept_NUC 1.00 (0.70) 1.00 (0.81) 1.00 (0.90)
mcg_ME2 0.90 (3.02) 0.61 (5.43) 0.63 (3.86)
mcg_ME3 0.92 (1.67) 0.82 (2.18) 0.92 (2.43)
mcg_MIT 0.90 (0.90) 0.90 (1.13) 0.98 (1.28)
mcg_NUC 0.96 (0.67) 0.94 (0.90) 0.98 (0.96)
gvh_ME2 0.94 (2.46) 0.56 (7.29) 0.83 (4.09)
gvh_ME3 0.94 (1.42) 0.82 (2.24) 0.94 (2.28)
gvh_MIT 0.92 (0.96) 0.86 (1.25) 0.95 (1.34)
gvh_NUC 0.94 (0.61) 0.93 (0.86) 0.96 (0.94)
alm_ME2 0.93 (3.53) 0.47 (12.2) 0.72 (4.91)
alm_ME3 0.95 (3.20) 0.56 (5.97) 0.62 (4.24)
alm_MIT 0.85 (1.22) 0.89 (1.50) 0.91 (1.50)
alm_NUC 0.93 (0.82) 0.90 (0.99) 0.98 (1.13)
mit_ME2 0.93 (2.50) 0.57 (5.91) 0.89 (4.13)
mit_ME3 0.97 (1.99) 0.84 (2.76) 0.92 (3.07)
mit_MIT 0.87 (0.93) 0.84 (1.10) 0.90 (1.09)
mit_NUC 0.93 (0.77) 0.90 (0.95) 0.94 (1.03)
vac_ME2 0.99 (2.07) 0.63 (3.67) 0.88 (3.66)
vac_ME3 0.97 (1.63) 0.87 (2.35) 0.94 (2.80)
vac_MIT 0.98 (0.76) 0.90 (1.08) 0.97 (1.14)
vac_NUC 0.95 (0.54) 0.89 (0.75) 0.94 (0.83)
nuc_ME2 0.98 (3.40) 0.55 (7.57) 0.94 (6.30)
nuc_ME3 0.74 (1.79) 0.74 (2.42) 0.85 (2.68)
nuc_MIT 0.94 (1.35) 0.83 (1.69) 0.95 (1.87)
nuc_NUC 0.89 (0.88) 0.81 (1.09) 0.81 (0.99)

### v2

@@ Appendix H Coverage rate of the marginal credible interval
Table 7: Coverage (rate) and median width (size) over 100 repetitions of marginal credible intervals for logistic regression with synthetic data. The method achieving at least the target coverage with the smallest size is shown in bold. If none of the methods achieve the target coverage, the method with the highest coverage is shown in bold. The target coverage is 0.95.
TabMGP $B$ Copula Bayes
Setup Feature Name Rate Size Rate Size Rate Size Rate Size
intercept_1.0 0.96 1.11 0.88 1.36 0.98 1.67 0.84 0.93
x10_1.0 0.86 1.76 0.66 2.07 0.99 2.24 0.69 1.25
x1_1.0 0.97 0.97 0.86 1.44 0.99 1.69 0.77 0.95
x2_1.0 0.85 1.19 0.76 1.56 0.98 1.82 0.71 1.01
x3_1.0 0.95 1.04 0.81 1.41 1.00 1.69 0.80 0.97
x4_1.0 0.98 0.99 0.85 1.41 0.99 1.71 0.73 0.94
x5_1.0 0.94 1.04 0.86 1.44 0.99 1.73 0.81 0.95
x6_1.0 0.90 1.34 0.77 1.56 1.00 1.88 0.75 1.04
x7_1.0 0.89 1.21 0.79 1.54 1.00 1.85 0.72 1.00
x8_1.0 0.85 1.48 0.75 1.74 0.99 2.01 0.70 1.11
Logistic x9_1.0 0.89 1.44 0.73 1.67 0.99 1.94 0.69 1.08
intercept_1.0 0.90 1.48 0.73 2.02 0.99 2.16 0.70 1.24
x10_1.0 0.91 2.53 0.52 3.37 1.00 3.00 0.56 1.87
x1_1.0 0.96 1.21 0.82 1.84 1.00 1.98 0.74 1.14
x2_1.0 0.86 1.61 0.70 2.09 0.97 2.16 0.71 1.27
x3_1.0 0.91 1.33 0.79 1.79 0.99 1.96 0.75 1.16
x4_1.0 0.97 1.26 0.79 1.77 0.99 2.00 0.73 1.16
x5_1.0 0.94 1.24 0.84 1.76 0.97 1.89 0.79 1.16
x6_1.0 0.91 1.86 0.68 2.37 1.00 2.37 0.67 1.35
x7_1.0 0.93 1.60 0.74 2.02 0.99 2.11 0.75 1.26
x8_1.0 0.90 2.07 0.62 2.76 0.99 2.51 0.63 1.50
GMM(0)(0) x9_1.0 0.95 1.92 0.64 2.54 1.00 2.33 0.66 1.43
intercept_1.0 0.96 1.10 0.89 1.36 0.98 1.67 0.80 0.94
x10_1.0 0.90 1.78 0.66 1.96 1.00 2.25 0.71 1.26
x1_1.0 0.96 0.96 0.82 1.46 1.00 1.68 0.76 0.96
x2_1.0 0.85 1.24 0.77 1.54 0.99 1.80 0.76 1.02
x3_1.0 0.91 1.06 0.83 1.43 0.99 1.68 0.78 0.97
x4_1.0 0.96 0.99 0.82 1.39 0.99 1.69 0.74 0.95
x5_1.0 0.93 1.02 0.88 1.42 1.00 1.71 0.79 0.95
x6_1.0 0.88 1.36 0.75 1.59 1.00 1.86 0.72 1.04
x7_1.0 0.86 1.18 0.78 1.55 1.00 1.82 0.74 0.98
x8_1.0 0.86 1.51 0.68 1.81 0.99 2.02 0.66 1.11
GMM(1)(-1) x9_1.0 0.94 1.46 0.73 1.65 1.00 1.92 0.66 1.07
intercept_1.0 0.92 1.05 0.84 1.35 0.98 1.67 0.79 0.89
x10_1.0 0.85 1.42 0.73 1.48 1.00 1.82 0.77 1.03
x1_1.0 0.97 0.81 0.89 1.22 0.97 1.56 0.86 0.85
x2_1.0 0.87 0.94 0.78 1.28 0.99 1.60 0.77 0.89
x3_1.0 0.94 0.85 0.88 1.24 0.97 1.58 0.78 0.86
x4_1.0 0.99 0.84 0.91 1.21 1.00 1.54 0.78 0.85
x5_1.0 0.90 0.88 0.82 1.26 1.00 1.53 0.78 0.86
x6_1.0 0.91 1.05 0.84 1.33 1.00 1.67 0.77 0.91
x7_1.0 0.83 0.97 0.84 1.33 1.00 1.61 0.77 0.87
x8_1.0 0.81 1.18 0.78 1.37 1.00 1.74 0.74 0.92
GMM(2)(-2) x9_1.0 0.83 1.08 0.79 1.33 1.00 1.65 0.71 0.91
@@
Table 8: Coverage (rate) and median width (size) over 100 repetitions of marginal credible intervals for linear regression with real data. The method achieving at least the target coverage with the smallest size is shown in bold. If none of the methods achieve the target coverage, the method with the highest coverage is shown in bold. The target coverage is 0.95.
TabMGP $B$ Copula Bayes
Setup Feature Name Rate Size Rate Size Rate Size Rate Size
age 0.83 0.36 0.83 0.35 0.83 0.32 0.81 0.29
cement 0.90 0.32 0.91 0.30 0.98 0.50 0.92 0.34
coarse_aggregate 0.91 0.29 0.91 0.27 0.99 0.47 0.94 0.30
fine_aggregate 0.95 0.29 0.95 0.26 1.00 0.45 0.96 0.31
fly_ash 0.96 0.34 0.92 0.32 0.99 0.63 0.96 0.36
intercept 0.91 0.29 0.92 0.26 1.00 0.55 0.95 0.28
concrete superplasticizer 0.91 0.37 0.85 0.35 0.99 0.49 0.90 0.36
col_1 1.00 0.23 0.88 0.44 0.99 0.50 0.95 0.45
col_2 0.93 0.25 0.94 0.51 1.00 0.80 0.88 0.41
col_3 1.00 0.24 0.95 0.50 1.00 0.81 0.91 0.40
quake intercept 0.96 0.47 0.95 0.50 1.00 1.02 0.96 0.39
chord_length 0.96 0.44 0.91 0.38 0.98 0.55 0.96 0.42
displacement_thickness 0.93 0.45 0.87 0.40 0.96 0.59 0.94 0.42
free_stream_velocity 0.94 0.40 0.93 0.35 0.99 0.55 0.95 0.41
frequency 0.90 0.48 0.87 0.43 0.84 0.47 0.88 0.42
airfoil intercept 0.93 0.42 0.93 0.36 0.99 0.63 0.95 0.40
glazing_area 0.93 0.25 0.86 0.15 1.00 0.34 1.00 0.42
glazing_area_distribution 0.96 0.25 0.89 0.15 0.99 0.33 1.00 0.43
intercept 0.92 0.28 0.88 0.15 1.00 0.45 1.00 0.40
orientation 0.96 0.24 0.88 0.15 1.00 0.34 1.00 0.41
overall_height 0.97 0.26 0.89 0.15 1.00 0.49 1.00 0.42
energy wall_area 1.00 0.25 0.96 0.12 1.00 0.35 1.00 0.43
CIC0 0.89 0.46 0.86 0.43 0.86 0.49 0.90 0.43
GATS1i 0.92 0.41 0.92 0.34 0.96 0.46 0.97 0.42
NdsCH 0.95 0.45 0.81 0.36 0.95 0.45 0.93 0.45
NdssC 0.94 0.46 0.89 0.44 0.94 0.47 0.93 0.48
SM1_Dz 0.89 0.48 0.83 0.38 0.92 0.48 0.88 0.45
fish intercept 0.93 0.43 0.83 0.35 0.99 0.54 0.90 0.41
intercept 0.94 0.47 0.91 0.38 0.90 0.35 0.94 0.42
theta1 0.96 0.41 0.85 0.38 0.86 0.36 0.92 0.44
theta2 0.98 0.40 0.86 0.37 0.88 0.35 0.96 0.43
theta3 0.96 0.52 0.90 0.36 0.78 0.34 0.94 0.43
theta4 0.96 0.42 0.90 0.39 0.90 0.34 0.93 0.43
theta5 0.82 0.44 0.87 0.40 0.86 0.33 0.91 0.42
theta6 0.92 0.43 0.79 0.39 0.78 0.34 0.85 0.43
theta7 0.94 0.42 0.84 0.38 0.84 0.35 0.91 0.43
kin8nm theta8 0.96 0.41 0.84 0.37 0.87 0.35 0.92 0.42
@@
Table 9: Coverage (rate) and median width (size) over 100 repetitions of marginal credible intervals for linear regression with real data. The method achieving at least the target coverage with the smallest size is shown in bold. If none of the methods achieve the target coverage, the method with the highest coverage is shown in bold. The target coverage is 0.95.
TabMGP $B$ Copula Bayes
Setup Feature Name Rate Size Rate Size Rate Size Rate Size
intercept 0.93 0.74 0.89 0.53 1.00 1.70 0.99 0.87
process.b1.capacity 0.98 0.50 0.94 0.40 1.00 1.04 0.99 0.51
process.b2.capacity 0.90 0.39 0.80 0.32 1.00 0.84 0.96 0.46
process.b3.capacity 0.92 0.33 0.94 0.26 0.98 0.66 1.00 0.48
process.b4.capacity 0.93 0.40 0.86 0.32 1.00 1.20 0.99 0.45
property.price 0.91 0.45 0.81 0.37 1.00 1.03 0.99 0.52
property.product_2 0.91 1.39 0.87 1.08 0.84 1.61 0.94 1.26
property.product_3 0.93 1.06 0.89 0.85 0.80 1.61 0.99 1.48
property.product_4 0.96 1.04 0.89 0.84 0.88 1.39 1.00 1.44
property.product_5 0.94 1.14 0.89 0.89 0.97 1.67 1.00 1.64
auction property.product_6 0.94 0.99 0.93 0.77 0.82 1.50 0.99 1.34
g1 0.92 0.39 0.82 0.30 0.86 0.34 0.97 0.45
g2 0.98 0.40 0.88 0.30 0.94 0.33 1.00 0.45
g3 0.96 0.39 0.86 0.30 0.89 0.33 0.97 0.45
g4 0.98 0.40 0.92 0.31 0.93 0.33 0.97 0.45
intercept 0.92 0.40 0.91 0.30 0.86 0.33 0.98 0.44
p2 0.97 0.36 0.90 0.29 0.96 0.34 0.98 0.45
p3 0.99 0.36 0.91 0.29 0.97 0.34 0.99 0.44
p4 0.99 0.38 0.87 0.29 0.97 0.34 0.98 0.44
tau1 0.91 0.39 0.81 0.29 0.91 0.32 0.96 0.44
tau2 0.94 0.39 0.89 0.29 0.89 0.33 0.99 0.44
tau3 0.91 0.38 0.86 0.29 0.90 0.33 0.96 0.44
grid tau4 0.94 0.40 0.89 0.30 0.91 0.33 0.98 0.44
Height 0.99 1.53 0.65 1.23 0.96 1.36 0.78 1.49
Sex_1 0.92 1.74 0.78 1.53 0.95 1.78 0.92 1.92
Sex_2 0.99 1.48 0.81 1.47 0.95 1.55 0.91 1.63
Shucked_weight 0.96 1.40 0.68 1.23 0.92 1.48 0.82 1.36
abalone intercept 0.94 1.23 0.76 1.01 0.99 1.76 0.91 1.23
@@
Table 10: Coverage (rate) and median width (size) over 100 repetitions of marginal credible intervals for logistic regression with real data. The method achieving at least the target coverage with the smallest size is shown in bold. If none of the methods achieve the target coverage, the method with the highest coverage is shown in bold. The target coverage is 0.95.
TabMGP $B$ Copula Bayes
Setup Feature Name Rate Size Rate Size Rate Size Rate Size
Eccentricity_Osmancik 0.91 5.36 0.37 7.52 0.98 14.46 0.62 4.81
Extent_Osmancik 0.99 1.19 0.93 2.11 1.00 4.14 0.93 1.40
Minor_Axis_Length_Osmancik 0.93 3.91 0.49 5.55 0.97 11.15 0.65 3.38
rice intercept_Osmancik 0.89 2.65 0.30 3.42 0.97 6.20 0.39 2.28
age_years_1 0.85 1.72 0.96 2.67 1.00 4.88 0.93 2.10
episode_number_1 1.00 1.19 0.93 2.54 1.00 3.28 0.99 1.83
intercept_1 1.00 1.77 0.98 2.91 1.00 6.42 0.97 1.95
sepsis sex_0male_1female_1_1 1.00 1.73 0.99 3.59 1.00 7.84 1.00 2.52
V1_2 0.96 2.24 0.96 2.62 1.00 2.85 0.98 1.96
V2_2 0.97 2.23 0.98 2.49 0.97 2.44 0.94 1.71
V4_2 0.98 1.99 0.98 2.25 0.99 2.30 0.92 1.38
banknote intercept_2 1.00 1.33 1.00 1.44 1.00 1.45 1.00 1.01
end_1 0.98 0.99 0.97 1.10 1.00 2.08 0.99 0.74
event_1 1.00 0.78 1.00 0.92 1.00 2.05 1.00 0.68
id_1 0.99 0.77 1.00 0.91 1.00 1.83 0.95 0.66
intercept_1 0.94 1.44 0.77 1.76 1.00 2.40 0.75 1.04
mozilla size_1 0.93 3.96 0.74 4.69 0.99 4.41 0.57 2.72
V1_2 1.00 3.00 1.00 3.29 1.00 3.23 0.99 1.59
V3_2 0.99 2.91 0.99 3.18 1.00 3.37 0.98 1.76
skin intercept_2 1.00 1.70 1.00 1.85 1.00 2.11 1.00 1.76
V1_2 0.82 1.80 0.91 2.64 0.98 4.18 0.83 1.63
V2_2 0.91 2.30 0.74 3.51 0.99 4.98 0.68 2.19
V4_2 0.81 1.76 0.79 2.74 1.00 4.01 0.77 1.91
blood intercept_2 1.00 1.48 0.93 2.14 1.00 3.68 0.92 1.28
V1_2 0.96 1.93 0.85 2.05 1.00 3.58 0.86 1.72
V2_2 0.95 1.84 0.87 2.15 0.98 3.44 0.80 1.46
V3_2 0.90 1.87 0.80 2.34 0.96 3.70 0.71 1.43
V4_2 0.91 1.68 0.80 1.97 0.95 3.68 0.75 1.24
V5_2 0.89 1.42 0.83 1.90 0.94 3.30 0.73 1.11
phoneme intercept_2 0.97 1.73 0.83 1.94 1.00 3.37 0.89 1.39
fAlpha_h 0.87 2.48 0.79 2.96 0.93 8.26 0.71 1.54
fAsym_h 0.96 2.55 0.78 3.38 1.00 7.77 0.72 2.26
fConc1_h 0.95 2.59 0.76 3.47 0.99 9.79 0.65 1.97
fDist_h 0.93 1.99 0.81 2.66 0.98 6.56 0.72 1.60
fLength_h 0.87 4.56 0.69 5.94 0.98 13.39 0.62 3.57
fM3Long_h 0.95 2.44 0.87 3.45 0.99 7.24 0.80 2.22
fM3Trans_h 0.97 2.21 0.88 3.12 1.00 6.63 0.81 1.92
fWidth_h 0.95 4.38 0.70 5.57 0.98 11.94 0.67 3.62
telescope intercept_h 0.97 1.93 0.88 2.54 1.00 7.44 0.83 1.56